OpenStudy (anonymous):

Three research departments have 10,9, and 7 members, respectively. Each department is to select a delegate and an alternate to represent the department at a conference. In how many ways can this be done?

OpenStudy (anonymous):

\[10\times 9\times 9\times 8\times 7\times 6\] hope method is clear

OpenStudy (kropot72):

10P2 + 9P2 + 7P2